# Catastrophic Liability: Managing Systemic Risks in Frontier AI Development

### v1

@@ Abstract
As artificial intelligence systems grow more capable and autonomous, frontier AI development poses potential systemic risks that could affect society at a massive scale. Current practices at many AI labs developing these systems lack sufficient transparency around safety measures, testing procedures, and governance structures. This opacity makes it challenging to verify safety claims or establish appropriate liability when harm occurs. Drawing on liability frameworks from nuclear energy, aviation software, and healthcare, we propose a comprehensive approach to safety documentation and accountability in frontier AI development.
AI Governance, Legal liability, AI Safety, Documentation, Auditing, TAIS
@@ 1 Introduction
As the capabilities of artificial intelligence (AI) systems rapidly advance, it is increasingly clear that frontier AI development has the potential to pose systemic risks to humanity if it is not done safely and responsibly. Policymakers, researchers, and industry leaders recognize the need to prioritize establishing robust frameworks for AI safety and governance (Center for AI Safety, 2023).
@@
The risks from AI emerge during development, not just adoption; if an advanced AI system escapes control to pursue its own goals, or is stolen and misused to accelerate chemical, biological, or nuclear weapons development, the damages could be catastrophic (Slattery et al., 2024). Yet the practices of companies developing frontier AI systems are often highly opaque, with little transparency into their safety practices, testing procedures, or governance structures for managing risks. This leaves the critical task of verifying the safety of these immensely powerful AI systems largely in the hands of the companies themselves or third-party evaluators contracted to test the final products.
@@
The resulting lack of oversight is deeply concerning given the stakes involved. Furthermore, competition dynamics incentivize AI companies to race ahead in capabilities at the expense of thorough safety practices (Walter, 2024; Bengio et al., 2024). Thus, policymakers and researchers have an interest in assessing whether self-reported safety claims truly mitigate key risks, but the opaque nature of AI development makes this difficult or impossible.
Contributions.
To address these challenges, we propose a new policy framework focusing on audits and transparency to ensure appropriate safety practices are followed in the development of frontier AI systems. Drawing on lessons from other high-risk domains, we argue that AI labs have a duty of care to implement, document and disclose responsible AI development practices to an independent auditor. Our analysis shows that effective governance requires:
  1. 1.
    Mandatory documentation of safety practices throughout the AI development lifecycle, not just at deployment;
  2. 2.
    Clear standards for demonstrating meaningful human oversight of increasingly capable systems
  3. 3.
    Independent third-party audit mechanisms that can assess novel risks while protecting trade secrets
  4. 4.
    Liability frameworks that establish responsibility across complex development pipelines
In the sections that follow, we first define key concepts like the duty of care and systemic risks from AI. We then examine case studies in nuclear energy, aviation and healthcare to identify relevant best practices and liability frameworks. Finally, we integrate these insights with existing policy discussions around the EU AI Act and emerging U.S. standards to propose feasible recommendations (see Figure 1).
Our goal is to ensure the transformative potential of artificial intelligence is steered in a positive direction by closing the gap between AI development practices and safety measures proportional to AI’s capabilities.
@@ 2.1 Duty of Care
The definition of ‘duty of care’ varies by jurisdiction. A general identification of ‘duty of care’ made by The Court of Justice of the European Union in their rulings is as an “information-related concept of decision-making accountability” (Hofmann, 2020). Across different rulings, there is a balance to be struck between the proportionality and discretion of the judicial board. Such a balance is not only important for respecting legislative bodies, but also for the discretion of different areas of expertise, such as boards of experts.
Though the proposed EU AI Liability Directive was abandoned in early 2025 (Andrews, 2025), its more specific analysis remains informative. According to Article 2(9): “‘duty of care’ means a required standard of conduct, set by national or Union law, in order to avoid damage to legal interests recognized at national or Union law level, including life, physical integrity, property and the protection of fundamental rights.”
Duty of care evaluations in the US often rely on “customs and standards.” (Abraham, 2009) However, precedent has also been set that customs cannot trump “reasonable practice.” For example, in 1932 a tugboat was said to breach duty of care by not including a functioning radio receiver, despite the fact that not including a radio receiver was common practice. The justification in this case was based on the relative cost of a functioning radio receiver versus the cost of the cargo lost due to lack of receiver (Epstein, 1992).
@@ 2.2 High Risk and Systemic Risk
The EU AI Act’s definitions of high and systemic-risk AI are important for understanding how EU courts will approach liability. Although other jurisdictions will not depend on EU precedent, we will analyze both EU and US guidance in Section 5. We review the types of risk considered by the EU to motivate our analogies to the liability precedent for similar risks. The EU AI Act makes an important distinction between “high-risk” AI systems and “general-purpose AI models with systemic risk.” High-risk AI systems are defined primarily by their intended uses and predictable misuses, particularly in areas e.g biometric identification, critical infrastructure, education, employment, essential services, law enforcement, migration, and justice administration. The key factor is whether misuse or malfunction of the AI system could threaten health, safety, or fundamental rights in these specific domains (European Parliament, 2024).
@@
The key distinction between these categories lies in their scope and timing. High-risk designations focus on specific applications and use cases, while systemic risk addresses underlying model capabilities that could create widespread impacts before any specific deployment. An AI system evaluating job candidates would be high-risk, but the foundation model on which it is built on may pose systemic risks during training itself, as it develops capabilities that could enable deception or circumvention of safety measures (Meinke et al., 2025).
@@ 3 Case Studies
To inform how liability should be treated in high-risk emerging technology like AI development, we examine other domains with established liability frameworks that share key characteristics with frontier AI: the potential for catastrophic harm, complex technical systems requiring specialized expertise, and challenges in balancing innovation with safety. We focus on nuclear energy, aviation software, and healthcare to understand how liability frameworks have evolved to address safety risks through transparency requirements and independent oversight.
Each case study illuminates different aspects relevant to the liability of AI development. Nuclear energy demonstrates how to handle rare but catastrophic risks through organizational governance requirements. Aviation software shows how complex autonomous systems can be made safer through rigorous documentation and third-party validation. Healthcare illustrates the challenges of attributing liability when AI systems are integrated into human decision-making processes. Together, these examples provide insights for developing appropriate liability frameworks for frontier AI development.
@@ 3.1 Nuclear Energy
It is critical to define the “who,” “when,” and “how much” of liability for harms we can anticipate. In the nuclear sector, catastrophic events such as radioactive material release could cause damages that exceed the industry’s ability to compensate. Similar concerns exist for AI development. A misaligned AI system with access to critical infrastructure could cause widespread disruption before adequate interventions are possible. Rather than waiting until such harms materialize, the discussion of liability must begin when risks are visible (Decker & Rauhut, 2021).
In the U.S., nuclear power accident liability for radiological damage is capped by the Price-Anderson Act of 1957 at the maximum insurance available to the nuclear energy industry. While this cap ensures victims would be compensated quickly if disaster strikes, it highlights the gap between possible compensation and potential catastrophic damages - the cap on liability does not limit the actual risk of harm (Holt, 2024).
@@
The nuclear industry illustrates both successful and problematic approaches to organizational accountability and governance as key drivers of risk management. Japan’s experience, as analyzed by Juraku & Sugawara, demonstrates how governance documentation that focuses on superficial compliance rather than active engagement with safety challenges can mask fundamental issues. The authors found that the Japanese nuclear industry’s tendency toward “strategic overlooking of ‘uncomfortable knowledge”’ and “automated decision making” contributed to the Fukushima disaster, despite the country’s formal governance procedures for nuclear risk management.
In contrast, effective governance frameworks drive internal practices and norms that facilitate a robust organizational risk culture. As found by Decker & Rauhut, strong governance can:
@@
    Drive internal practices and norms that facilitate organizational risk culture;
  2. 2.
    Help document security practices without compromising sensitive information;
  3. 3.
    Enable assessment of duty of care based on industry best practices, responses to known threats, and continuous improvement;
  4. 4.
    Support organizational accountability in preventative risk management.
Courts assess nuclear operators’ duty of care based not only on regulatory compliance but also on broader standards including industry best practices, responses to known threats, implementation of audit recommendations, and organizational capacity for continuous improvement. Governance documentation in the nuclear industry covers risk management communication at executive levels, security culture development, and risk mitigation strategies (Decker & Rauhut, 2021).
Drawing together these lessons, we argue that the NIST AI Risk Management Framework provides a strong foundation for standardizing AI risk management practices, but it needs strengthening in several areas. Its next iteration should:
  1. 1.
    Require documentation showing active engagement with safety challenges, not just compliance;
@@
    Mandate processes that prevent over-reliance on automated decision-making about risks;
@@
The parallel with nuclear safety is particularly relevant since both industries deal with the risks of rare but catastrophic failures that could exceed any possible compensation. This suggests the need for liability frameworks that focus on preventing harm through documented safety practices rather than just compensating after incidents occur.
We suggest that following the lead of the Price-Anderson Act of 1957, these practices can be enforced via certification by a federal body. Although there would be advantages to the enforcement capabilities of requiring providers of risky models to obtain certification in order to operate, we anticipate that industry push back would make this politically untenable. However, extending the Price-Anderson Act’s insurance-based liability cap for certified operators could make this feasible.
@@ 3.2 Aviation Software
Negligence in aircraft software can have catastrophic consequences, including loss of life and environmental damage. However, lessons from aviation’s long history have led to automation standards, rigorous training, and robust regulations. Aviation software is subject to strict compliance standards such as DO-178C, which mandate documentation and traceability throughout the software life cycle. (Federal Aviation Administration, 2021) This transparency enables audits by authorities such as the FAA and EASA, ensuring adherence to safety regulations, with legal consequences for non-compliance. Aviation accidents can result from breaches of duty of care, including human, technical, and organizational failures, each with distinct liability.
@@
Artificial intelligence is increasingly used in aviation: predictive maintenance, aircraft manufacturing, automating and enhancing air traffic management, crew management, messaging automation. (Symphony Solutions, 2024) Regulatory frameworks are emerging to address AI risks. The EASA AI Roadmap 2.0 emphasizes a risk-based certification process, ensuring that AI is understandable and allows for human intervention. (European Union Aviation Safety Agency, 2023) The FAA’s AI roadmap establishes liability and transparency requirements, treating AI as a tool rather than an autonomous entity (Federal Aviation Administration, 2023); it mandates clear distinctions between human and automated responsibilities and rigorous testing for safety compliance. Although these roadmaps are positive steps, it is unclear whether current liability measures are sufficient to address catastrophic AI failures. Both agencies must explicitly tackle transparency concerns, over-reliance on AI, and decision-making risks.
@@
Lessons from aviation software liability frameworks can inform AI governance. First, aviation’s comprehensive documentation and third-party validation requirements (e.g., DO-178C) provide a model for AI transparency. Second, aviation treats automation as a tool with human accountability, making it clear where liability falls in the value chain from development to application. Third, aviation imposes stringent regulations on high-risk systems, illustrating a risk-sensitive regulatory model that could usefully inform AI oversight.
However, AI development presents unique challenges. Unlike aviation software, AI may gain unforeseen capabilities, pose risks during the development process itself, or be applied unpredictably. This calls for new liability frameworks that:
  1. 1.
    Mandate comprehensive documentation of safety measures throughout the AI development process, not just at deployment;
  2. 2.
    Establish clear responsibility and testing requirements for general-purpose AI capabilities;
  3. 3.
    Create robust third-party audit mechanisms that can assess novel risks;
  4. 4.
    Develop standards for demonstrating meaningful human oversight of increasingly capable AI systems.
@@
The evolution of aviation safety standards which is shaped by the lessons of past failures, underscores the importance of establishing strong liability frameworks for AI systems. Although the specific technical standards used in aviation may not apply directly to AI, the broader principles of documentation, responsibility, and independent validation provide important guidance for developing effective AI governance.

### v2

@@ Abstract
As artificial intelligence systems grow more capable and autonomous, frontier AI development poses potential systemic risks that could affect society at a massive scale. Current practices at many AI labs developing these systems lack sufficient transparency around safety measures, testing procedures, and governance structures. This opacity makes it challenging to verify safety claims or establish appropriate liability when harm occurs. Drawing on liability frameworks from nuclear energy, aviation software, cybersecurity, and healthcare, we propose a comprehensive approach to safety documentation and accountability in frontier AI development.
@@ 1 Introduction
As the capabilities of artificial intelligence (AI) systems rapidly advance, it is increasingly clear that frontier AI development has the potential to pose systemic risks to humanity if it is not done safely and responsibly. Policymakers, researchers, and industry leaders recognize the need to prioritize establishing robust frameworks for AI safety and governance (Center for AI Safety 2023).
@@
The risks from AI emerge during development, not just adoption; if an advanced AI system escapes control to pursue its own goals, or is stolen and misused to accelerate chemical, biological, or nuclear weapons development, the damages could be catastrophic (Slattery et al. 2024). Yet the practices of companies developing frontier AI systems are often highly opaque, with little transparency into their safety practices, testing procedures, or governance structures for managing risks. This leaves the critical task of verifying the safety of these immensely powerful AI systems largely in the hands of the companies themselves or third-party evaluators contracted to test the final products.
@@
The resulting lack of oversight is deeply concerning given the stakes involved. Furthermore, competition dynamics incentivize AI companies to race ahead in capabilities at the expense of thorough safety practices (Walter 2024; Bengio et al. 2024). Thus, policymakers and researchers have an interest in assessing whether self-reported safety claims truly mitigate key risks, but the opaque nature of AI development makes this difficult or impossible.
Thesis and approach.
We argue that greater transparency and documentation of risk management practices can simultaneously increase safety and reduce costs for frontier AI development. Each of the major stakeholders involved—policymakers, standards bodies, and industry—can advance their own interests while supporting each other by describing and demonstrating responsibility. This mutually beneficial dynamic emerges from three key contributions:
(1) We establish that voluntary AI standards already carry legal force through existing tort law. U.S. courts determine negligence based on ”reasonable practice” rather than merely common practice, giving frameworks like NIST’s AI Risk Management Framework potential legal weight even without mandatory regulation. This challenges the assumption that voluntary commitments lack enforcement mechanisms.
(2) We demonstrate that frontier AI already poses systemic risks requiring expanded duties of care. Unlike high-risk AI applications that create dangers through specific deployments, general-purpose models with systemic risk can pose threats through their underlying capabilities—regardless of intended use. The EU AI Act recognizes this distinction by imposing obligations on frontier model providers based on computational thresholds and capability assessments, not just deployment contexts.
(3) We show how comprehensive documentation creates competitive advantages while improving safety. Through case studies from nuclear energy, aviation, healthcare, and cybersecurity, we demonstrate that transparency reduces liability exposure for responsible developers while enabling meaningful oversight. Current opacity leaves all parties worse off.
@@
Structure and contributions.
To substantiate these arguments, we first establish the legal foundations of duty of care and distinguish between high-risk AI applications and systemic risks from frontier models (Section 2). We then examine four industries that have regularly confront safety-critical risks: nuclear energy’s governance frameworks that prevent ”strategic overlooking” of risks; aviation’s rigorous third-party validation; healthcare’s attribution challenges in human-AI systems; and cybersecurity’s different transparency requirements for deployment versus development risks (Section 3). Drawing on current EU and U.S. regulatory developments (Section 4), we synthesize concrete recommendations showing how each stakeholder can benefit from comprehensive safety documentation (Section 5, also see Figure 1).
Our goal is to ensure the transformative potential of artificial intelligence is steered in a positive direction by closing the gap between AI development practices and safety measures proportional to AI’s capabilities—not through unrealistic regulatory enthusiasm, but by demonstrating how transparency serves everyone’s interests.
@@ 2.1 Duty of Care
The definition of ‘duty of care’ varies by jurisdiction. A general identification of ‘duty of care’ made by The Court of Justice of the European Union in their rulings is as an “information-related concept of decision-making accountability” (Hofmann 2020). Across different rulings, there is a balance to be struck between the proportionality and discretion of the judicial board. Such a balance is not only important for respecting legislative bodies, but also for the discretion of different areas of expertise, such as boards of experts.
Though the proposed EU AI Liability Directive was abandoned in early 2025 (Andrews 2025), its more specific analysis remains informative. According to Article 2(9): “‘duty of care’ means a required standard of conduct, set by national or Union law, in order to avoid damage to legal interests recognized at national or Union law level, including life, physical integrity, property and the protection of fundamental rights.”
Duty of care evaluations in the US often rely on “customs and standards.” (Abraham 2009) However, precedent has also been set that customs cannot trump “reasonable practice.” For example, in 1932 a tugboat was said to breach duty of care by not including a functioning radio receiver, despite the fact that not including a radio receiver was common practice. The justification in this case was based on the relative cost of a functioning radio receiver versus the cost of the cargo lost due to lack of receiver (Epstein 1992).
@@ 2.2 High Risk and Systemic Risk
The EU AI Act’s definitions of high and systemic-risk AI are important for understanding how EU courts will approach liability. Although other jurisdictions will not depend on EU precedent, we will analyze both EU and US guidance in Section 5. We review the types of risk considered by the EU to motivate our analogies to the liability precedent for similar risks. The EU AI Act makes an important distinction between “high-risk” AI systems and “general-purpose AI models with systemic risk.” High-risk AI systems are defined primarily by their intended uses and predictable misuses, particularly in areas e.g biometric identification, critical infrastructure, education, employment, essential services, law enforcement, migration, and justice administration. The key factor is whether misuse or malfunction of the AI system could threaten health, safety, or fundamental rights in these specific domains (European Parliament 2024).
@@
The key distinction between these categories lies in their scope and timing. High-risk designations focus on specific applications and use cases, while systemic risk addresses underlying model capabilities that could create widespread impacts before any specific deployment. An AI system evaluating job candidates would be high-risk, but the foundation model on which it is built on may pose systemic risks during training itself, as it develops capabilities that could enable deception or circumvention of safety measures (Meinke et al. 2025).
@@ 3 Case Studies
To inform how liability should be treated in high-risk emerging technology like AI development, we examine other domains with established liability frameworks that share key characteristics with frontier AI: the potential for catastrophic harm, complex technical systems requiring specialized expertise, and challenges in balancing innovation with safety. We focus on nuclear energy, aviation software, healthcare, and cybersecurity to understand how liability frameworks have evolved to address safety risks through transparency requirements and independent oversight.
Each case study illuminates different aspects relevant to the liability of AI development. Nuclear energy demonstrates how to handle rare but catastrophic risks through organizational governance requirements. Aviation software shows how complex autonomous systems can be made safer through rigorous documentation and third-party validation. Healthcare illustrates the challenges of attributing liability when AI systems are integrated into human decision-making processes. Cybersecurity highlights adversarial risk, anchoring liability in ongoing monitoring, coordinated vulnerability disclosure, and a developer duty to patch. Together, these examples provide insights for developing appropriate liability frameworks for frontier AI development.
Recent commentators disagree on how tightly analogies to high-risk industries should constrain liability design. What we call an integrationist strand adopts the view that enterprise-level control structures already mandated for safety-critical domains can be transposed to frontier AI with only minor adjustment (National Institute of Standards and Technology 2023; ISO 2023). An empirical approach, exemplified by Rismani and colleagues, shows that classic aviation and medical-device tools such as STPA and FMEA do surface socio-technical hazards in machine-learning pipelines, yet they must be supplemented with continuous post-deployment monitoring to remain workable (Rismani et al. 2023). A third, pluralist strand argues for matching each liability instrument to the class of harm it best addresses: Murray links adversarial misuse to finance-style “risk-owner” models, while Koessler and Schuett recommend customized toolkits assembled from finance, nuclear energy and biocontainment practice (Murray 2025; Koessler and Schuett 2023). The four case studies developed in this paper align naturally with this pluralist perspective.
@@ 3.1 Nuclear Energy
Among the industries we examine, nuclear energy provides the most direct parallel to frontier AI development in terms of catastrophic risk potential and the need for proactive liability frameworks. Both sectors face a critical challenge: defining the ”who,” ”when,” and ”how much” of liability for harms that could exceed any industry’s ability to compensate. In nuclear energy, radioactive material release represents this catastrophic potential; in AI development, a misaligned system with access to critical infrastructure could cause widespread disruption before adequate interventions are possible. Rather than waiting until such harms materialize, the discussion of liability must begin when risks are visible (Decker and Rauhut 2021).
Catastrophic mistake potential.
Nuclear power plant operation offers particularly relevant lessons for AI development. Both involve beneficial technologies with risks of catastrophic accidental harm, where failure modes could exceed traditional compensation mechanisms. While critics have rightly noted that nuclear weapons involve centralized control, clear physical constraints, and discrete catastrophic events that don’t apply to AI systems (Kapoor 2025), our focus on civilian power generation addresses these concerns. The nuclear power industry’s liability frameworks thus provide relevant lessons for managing systemic risks from technologies whose failure modes could exceed traditional compensation mechanisms—AI included (Goudarzi 2024; Matthews 2023).
In this spirit, our analogy is not between the arms races in AI and nuclear weapons. Instead, the disasters we compare in this case study are nuclear reactor meltdowns and misaligned AI takeover. We recognize that relative to nuclear energy, the sources and likelihood of such catastrophic risks from AI can seem unclear (Baum 2025). However, progress in AI risk ontologies (Slattery et al. 2024; Bagehorn et al. 2025), makes it easier than ever to translate the scenario-oriented regulatory tool of safety cases from nuclear energy (Bounds 2020) and other safety-critical industries (Leveson 2011) to AI (Buhl et al. 2024; Hilton et al. 2025; Habli et al. 2025).
Beyond the scale of potential harm, the governance structures surrounding these technologies prove equally instructive.
Governance frameworks.
@@
The nuclear industry demonstrates how governance structures fundamentally shape risk management outcomes. The contrast between effective and problematic governance is stark. Juraku and Sugawara found that Japan’s pre-Fukushima nuclear governance encouraged ”automated decision making” and ”strategic overlooking of ’uncomfortable knowledge’,” allowing critical safety issues to persist beneath formal compliance procedures. This superficial approach to risk management contributed directly to the 2011 disaster.
Drawing on their multi-year study of nuclear security incentives, Decker and Rauhut found that effective governance extends beyond regulatory compliance to create organizational cultures that actively identify and mitigate risks. Their research revealed that market actors—including insurers, credit rating agencies, and courts—increasingly evaluate nuclear operators based on comprehensive governance practices rather than minimum regulatory adherence.
Decker and Rauhut’s Good Governance Template addresses these shortcomings by establishing principles that: (1) drive active engagement with emerging risks rather than checkbox compliance; (2) create transparent documentation processes that satisfy external stakeholders while protecting sensitive information; (3) demonstrate due care through continuous improvement and adoption of industry best practices; and (4) establish clear accountability for preventative risk management. Their research with judges, attorneys, and insurers confirmed that such comprehensive governance documentation can significantly reduce liability exposure and improve access to insurance and financing.
The lesson for AI governance is clear. Like nuclear technology, frontier AI requires governance frameworks that incentivize genuine engagement with emerging risks rather than superficial adherence to predetermined safety procedures.
The risks from inadequate governance become especially pronounced when considering how frontier AI companies deploy their systems. Recent analysis suggests that these companies may use their most advanced systems internally for extended periods to accelerate research and development, creating self-reinforcing cycles where AI systems help develop their successors with diminishing human oversight (Stix et al. 2025). Such internally deployed AI systems could operate with expanded permissions and potentially establish persistence beyond company systems (Pan et al. 2025) while remaining invisible to external regulators. Without robust governance frameworks like those developed for nuclear energy, these internal deployments could replicate the ”automated decision making” and ”strategic overlooking” patterns that preceded Fukushima.
These governance principles gain teeth through specific regulatory mechanisms.
Inspections and liability caps.
In the U.S., nuclear power accident liability for radiological damage is capped by the Price-Anderson Act of 1957 at the maximum insurance available to the nuclear energy industry. While this cap ensures victims would be compensated quickly if disaster strikes, it highlights the gap between possible compensation and potential catastrophic damages - the cap on liability does not limit the actual risk of harm (Holt 2024).
Chain of causation.
Recommendations.
Drawing together these lessons, we argue that the NIST AI Risk Management Framework provides a strong foundation for standardizing AI risk management practices, but it needs strengthening in several areas. Specifically, the framework’s next iteration should incorporate six key improvements drawn from nuclear safety experience:
@@
    Describe processes for documenting active engagement with safety challenges, not just compliance;
@@
    Define appropriate roles and processes to prevent over-reliance on automated decision-making about risks;
@@
We suggest that following the lead of the Price-Anderson Act of 1957, these practices can be enforced via certification by a federal body. While mandatory certification faces political challenges, the Price-Anderson model demonstrates how liability caps can incentivize voluntary adoption of rigorous safety standards—a particularly relevant approach for encouraging frontier AI labs to embrace comprehensive documentation requirements.
@@ 3.2 Aviation Software
Negligence in aircraft software can have catastrophic consequences, including loss of life and environmental damage. However, lessons from aviation’s long history have led to automation standards, rigorous training, and robust regulations. Aviation software is subject to strict compliance standards such as DO-178C, which mandate documentation and traceability throughout the software life cycle. (Federal Aviation Administration 2021) This transparency enables audits by authorities such as the FAA and EASA, ensuring adherence to safety regulations, with legal consequences for non-compliance. Aviation accidents can result from breaches of duty of care, including human, technical, and organizational failures, each with distinct liability.
@@
Artificial intelligence is increasingly used in aviation: predictive maintenance, aircraft manufacturing, automating and enhancing air traffic management, crew management, messaging automation. (Symphony Solutions 2024) Regulatory frameworks are emerging to address AI risks. The EASA AI Roadmap 2.0 emphasizes a risk-based certification process, ensuring that AI is understandable and allows for human intervention. (European Union Aviation Safety Agency 2023) The FAA’s AI roadmap establishes liability and transparency requirements, treating AI as a tool rather than an autonomous entity (Federal Aviation Administration 2023); it mandates clear distinctions between human and automated responsibilities and rigorous testing for safety compliance. Although these roadmaps are positive steps, it is unclear whether current liability measures are sufficient to address catastrophic AI failures. Both agencies must explicitly tackle transparency concerns, over-reliance on AI, and decision-making risks.
Lessons for AI Governance.
@@
Aviation’s software liability framework offers valuable insights for AI oversight. First, aviation’s comprehensive documentation and third-party validation requirements (e.g., DO-178C) provide a model for AI transparency. Importantly, these requirements have not stifled aviation innovation but rather enabled it by reducing liability uncertainty and facilitating meaningful oversight—supporting our thesis that transparency benefits all stakeholders. Second, aviation treats automation as a tool with human accountability, making it clear where liability falls in the value chain from development to application. This precedent for maintaining clear accountability chains remains relevant even as AI systems become increasingly autonomous. Third, aviation imposes stringent regulations on high-risk systems, illustrating a risk-sensitive regulatory model that could usefully inform AI oversight.
However, AI development presents unique challenges. Unlike aviation software, AI may gain unforeseen capabilities, pose risks during the development process itself, or be applied unpredictably. This distinction reinforces our argument about systemic risks from frontier AI models requiring expanded duty of care beyond deployment-specific risks. Aviation’s current struggle to adapt its validation methods to AI’s emergent capabilities highlights the need for new approaches that can assess not just intended functions but potential capabilities that may arise during development or deployment.
@@
The evolution of aviation safety standards, which is shaped by the lessons of past failures, underscores the importance of establishing strong liability frameworks for AI systems. Although the specific technical standards used in aviation may not apply directly to AI, the broader principles of documentation, responsibility, and independent validation provide important guidance for developing effective AI governance.